OpenStudy (anonymous):

The volume V of a cone is increasing at the rate of 28pi cubic units per second. At the instant when the radius r of the cone is 3 units, its volume is 12pi cubic units and the radius is increasing at 1/2 unit per second.

OpenStudy (anonymous):

At the instant when the radius of the cone is 3 units, what is the instantaneous rate of change of the area of its base with respect to its height???

OpenStudy (anonymous):

\[\frac{ dV }{ dt }=28\pi\] \[\frac{ dr }{ dt }=1/2\]

OpenStudy (anonymous):

\[V=\frac{ 1 }{ 3 }\pi r^2h\] \[V=12,r=3\]find h

OpenStudy (anonymous):

\[V=\frac{ 1 }{ 3 }Ah,A=\frac{ 3V }{ h }\] \[\frac{ dA }{ dt }=\frac{ -3V }{ h^2 }\] \[\frac{ -3(12) }{ h^2 }\] use the height above

OpenStudy (anonymous):

h=4?

OpenStudy (anonymous):

is there anther ? cos we didnot usethe rates above ,is there a next ?